Opened 10 years ago

Last modified 2 years ago

#245 new flight processing

GB09/00, flight day 154b/2009, Oxford Parks

Reported by: jatt Owned by: emca
Priority: whenever Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: peland Other processors: mark1

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB09_00-2009_154b_Oxford_parks

Data arrived from ARSF via SATA disk M on 9th June 2009.

Scientific objective: Calibration Target

Priority: None

PI: M Pollard (T Johnson is PhD working on this)

Sensors:

  • Eagle (16/12/2009)
  • Hawk (16/12/2009)
  • Leica LIDAR (17/05/2010)

Change History (20)

comment:1 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:2 Changed 10 years ago by jatt

Applanix has been processed and copied to ~arsf/workspace/GB09_00-2009_154b_Oxford_parks.

This directory is currently being used as a test for the part automation of the arsf processing.

comment:3 Changed 10 years ago by jatt

Hawk files have been renamed to the SWIRsh convention.

All sct offsets have been processed using the new automated approach and are the images are ready for selection.

comment:4 Changed 10 years ago by emca

  • Owner set to emca

comment:5 Changed 9 years ago by mggr

  • Description modified (diff)

Gary says to deliver to Prof. Mark Pollard. Tony Johnson is the PhD working on this.

comment:6 Changed 9 years ago by mggr

Approximate location of ground targets:

  • 451600E 206955N (BNG)
  • level 1 pixel coords (approx)
    • line 1: 238, 1753
    • line 2: 374, 1633

The targets are three 6m square (polysomething) cloth, organised on the rooftop of the building at the location above. Looking at it visually, the most northerly is white, then there's a ~2m gap, then a black target, then another gap (~6m? includes a walkway), then the gray target.

There is definite cloud corruption over the targets in all images.

comment:7 Changed 9 years ago by mggr

  • Cc peland added

Peter: see this ticket for locations of targets.

comment:8 Changed 9 years ago by mark1

Vectors have arrived.

comment:9 Changed 9 years ago by emca

Ready for delivery checking. Delivery directory created at
~airborne/workspace/GB09_00-2009_154b_Oxford_parks/delivery

comment:10 Changed 9 years ago by mark1

  • Other processors set to mark1

Delivery checking

comment:11 Changed 9 years ago by mark1

Updated data quality report to December version, minor corrections made to read_me file, all data looked at and example commands tested, results good.

Ready to deliver

comment:12 Changed 9 years ago by mark1

Photos have been processed and are ready to deliver.

comment:13 Changed 9 years ago by emca

LIDAR processing is complete - ready for delivery check at
~airborne/workspace/GB09_00-2009_154b_Oxford_parks/delivery/20100514/GB09_00

comment:14 Changed 9 years ago by mark1

LIDAR delivery has been checked and looks good.

Ready to deliver.

comment:15 Changed 9 years ago by emca

  • Description modified (diff)

LIDAR data sent to PI 17/05/2010

Hyperspectral and Photography sent 16/12/2009


comment:16 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been rsync'd to
~arsf/arsf_data/2009/flight_data/uk/GB09_00-2009_154b_Oxford_parks

Ready for archiving.

comment:17 Changed 6 years ago by stgo

SCTs from config file

Flightline Eagle Hawk
8 0.00 -0.08
9 -0.06 -0.10

comment:18 Changed 6 years ago by stgo

Copied raw header file band information from 132a. Original header files kept as copy, would not process due to calibration error.

Last edited 6 years ago by stgo (previous) (diff)

comment:19 Changed 2 years ago by asm

Archiving

Started. Both raw data and deliveries need archiving. Pushing data to the archive.

comment:20 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.